OpenStudy (anonymous):

120 factor tree 200 factor tree

OpenStudy (anonymous):

120=2*2*2*3*5 200=2*2*2*5*5

OpenStudy (anonymous):

1,2,3,4,5,6,8,

OpenStudy (chaise):

1,2,3,4,5,6,8,10,12,15,20,24,30,40,60,120.

OpenStudy (chaise):

1,2,4,5,8,10,20,25,40,50,100,200

OpenStudy (anonymous):

i meant that only

OpenStudy (anonymous):

i meant that only

OpenStudy (anonymous):

it's obvious from the prime factor multiplication..........

OpenStudy (shadowfiend):

If you're going to ask tons and tons of homework questions, chelsea4310, at least have the courtesy to give medals to the people who help you ;)